OpenStudy (anonymous):

Verify: 1/tanx - sinx/(cosx-1) = -cscx I've finished most of my homework but this problem is getting on my nerves! I just need to know how to solve it, not just the answer. It's probably easy, but my brain is fried...

OpenStudy (michele_laino):

hint: \[\frac{1}{{\tan x}} - \frac{{\sin x}}{{\cos x - 1}} = \frac{{\cos x}}{{\sin x}} - \frac{{\sin x}}{{\cos x - 1}}\]

OpenStudy (anonymous):

Well I got that part down, but I kept thinking it was wrong because for some reason I can't get passed that point.

OpenStudy (michele_laino):

we have to compute the least common multiple between sinx and cosx-1, which is: sinx(cosx-1)

OpenStudy (michele_laino):

\[\frac{{\cos x}}{{\sin x}} - \frac{{\sin x}}{{\cos x - 1}} = \frac{{\cos x\left( {\cos x - 1} \right) - \sin x\sin x}}{{\sin x\left( {\cos x - 1} \right)}} = ...?\]

OpenStudy (bee_see):